OpenStudy (anonymous):

Solve for x: 5|2x - 2| + 8 = 18. a:x = 2, x = -2 c:x = 2, x = 0 There are other choices, but I know its either a or c. Please HELP I will give a fan!

OpenStudy (anonymous):

@phi @SolomonZelman

OpenStudy (jdoe0001):

\(5|2x - 2| + 8 = 18\implies 5|2x - 2|=18-8\implies |2x - 2|=\cfrac{18-8}{5} \\ \quad \\ \large { |2x - 2|=\cfrac{18-8}{5}\to \begin{cases} +(2x - 2)=\cfrac{18-8}{5} \\ \quad \\ -(2x - 2)=\cfrac{18-8}{5} \end{cases} }\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 5|2x - 2| + 8 = 18 }\) \(\normalsize\color{blue}{ 5|2x - 2| + 8 \color{red}{-8}= 18\color{red}{-8} }\) \(\normalsize\color{blue}{ 5|2x - 2| = 10 }\) \(\normalsize\color{blue}{ 5|2x - 2| \color{red}{\div 5} = 10\color{red}{\div 5} }\) \(\normalsize\color{blue}{ |2x - 2| =2 }\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 2|x - 1| =2 }\) \(\normalsize\color{blue}{ 2|x - 1|\color{red}{ \div2} =2\color{red}{ \div2} }\) \(\normalsize\color{blue}{ |x - 1| =1 }\)

OpenStudy (solomonzelman):

can you finish it ?

OpenStudy (anonymous):

Yes I can thank you! :)

OpenStudy (anonymous):

both of you

OpenStudy (solomonzelman):

Anytime:)